# Beyond Diagonal RIS: Passive Maximum Ratio Transmission and Interference Nulling Enabler

### v1

@@ III-B Passive Interference Nulling
Such problem is non-convex and challenging to solve. Therefore, AO is used to find $\boldsymbol{\Theta}$ that complies with the previously mentioned constraints by projecting a certain initial solution into $\mathcal{S}_{\mathrm{null}}$, $\mathcal{S}_{\mathrm{loss}}$ and $\mathcal{S}_{\mathrm{recip}}$ alternatively. Fortunately, the symmetric and unitary projections have been presented with closed-form solutions in (11)–(12), while the interference nulling projection follows [10, Eq. (21a)] such that
$\Pi_{\mathcal{S}_{\mathrm{null}}}\left(\boldsymbol{\Theta}\right)=\mathrm{reshape}\left(\mathbf{v}\left(\boldsymbol{\Theta}\right),[N,N]\right)$ (21)
where $\mathbf{v}\left(\boldsymbol{\Theta}\right)\!=\!\mathrm{vec}\!\left(\boldsymbol{\Theta}\right)-\bar{\mathbf{A}}^{\ast}\!\left(\bar{\mathbf{A}}^{T}\bar{\mathbf{A}}^{\ast}\right)^{-1}\!\bar{\mathbf{A}}^{T}\mathrm{vec}\!\left(\boldsymbol{\Theta}\right)\!\in\mathbb{C}^{N^{2}\times 1}$.
Input: $\bar{\mathbf{A}},\,\mathrm{Iter},\,\epsilon,\,\varepsilon$
@@
1 Initialize $\boldsymbol{\Theta}^{(0)}$
2 $\tilde{\boldsymbol{\Theta}}^{(0)}=\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\Pi_{\mathcal{S}_{\mathrm{loss}}}\left(\boldsymbol{\Theta}^{(0)}\right)\right)$
3 for $i=:1:\mathrm{Iter}$ do
4       $\boldsymbol{\Theta}^{(i)}=\Pi_{\mathcal{S}_{\mathrm{null}}}\left(\tilde{\boldsymbol{\Theta}}^{(i-1)}\right)$
5       $\tilde{\boldsymbol{\Theta}}^{(i)}=\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\Pi_{\mathcal{S}_{\mathrm{loss}}}\left(\boldsymbol{\Theta}^{(i)}\right)\right)$
6
7      $\delta=\frac{\left\|\bar{\mathbf{A}}^{T}\mathrm{vec}\left({\tilde{\boldsymbol{\Theta}}^{(i)}}\right)\right\|^{2}_{2}-\left\|\bar{\mathbf{A}}^{T}\mathrm{vec}\left(\tilde{\boldsymbol{\Theta}}^{(i-1)}\right)\right\|^{2}_{2}}{\left\|\bar{\mathbf{A}}^{T}\mathrm{vec}\left(\tilde{\boldsymbol{\Theta}}^{(i-1)}\right)\right\|^{2}_{2}}$
8       if $\delta\leq\epsilon$ or ${\scriptstyle{\left\|\bar{\mathbf{A}}^{T}\mathrm{vec}\left(\tilde{\boldsymbol{\Theta}}^{(i)}\right)\right\|^{2}_{2}}}<\varepsilon$ or $i>\mathrm{Iter}$ then
9             break and Go to Line $9$
10
@@ 1 Initialize $\boldsymbol{\Theta}^{(0)}$
Figure 2: Convergence vs. number of iterations for interference nulling using BD-RIS and D-RIS with random initialization, where $K=8$ and $N=144$. a) Nulling condition $l^{2}$ norm. b) Nulling condition $l^{2}$ norm relative change.
@@
A pseudo-code that illustrates the steps to obtain $\boldsymbol{\Theta}^{\star}$ through AO is given in Algorithm 1. The algorithm needs the interference channels matrix $\bar{\mathbf{A}}$ as input. It starts by initializing the scattering matrix, which can be either a random initialization or a specific initialization. Furthermore, the initialized scattering matrix is projected to the lossless and reciprocal sets. Lines 33338888 illustrate the iterative AO which projects the $\boldsymbol{\Theta}$ from the previous iteration to interference nulling set and the lossless and reciprocal sets. This continues for a certain number of iterations until the breaking condition on line $7$ is met. On line $6$, $\delta$ is computed which is the $l^{2}$ norm for the nulling condition difference between the current iteration and the previous iteration. Finally, the symmetric unitary $\boldsymbol{\Theta}^{\star}$ that achieves interference nulling is returned. It is worth noting that $\epsilon$ and $\varepsilon$ determine the algorithm convergence speed, where the former is the relative rate of change tolerance and the latter is the nulling condition norm tolerance.
@@
Fig. 2 compares the convergence of Algorithm 1 with the D-RIS interference nulling AO algorithm [10]. Random initialization is assumed for both, with $N=144$, $K=8$, $\epsilon=10^{-6}$, $\varepsilon=10^{-10}$, $\mathrm{Iter}=10^{4}$, and $100$ realizations are considered for Monte-Carlo simulation. Also, i.i.d Rayleigh fading channels are considered. The curves represent the average convergence of the algorithm over $10^{4}$ iterations, while the shades represent the possible variations, i.e., the maximum and minimum convergence. It can be observed that BD-RIS achieves a rapid interference nulling, reaching $10^{-8}$ with significantly fewer iterations compared to D-RIS. In addition, the relative change for BD-RIS stabilizes after fewer iterations compared to D-RIS. Moreover, BD-RIS maintains a stable pattern, effectively achieving interference nulling, unlike D-RIS which is more dependent on the initialization of the scattering matrix, resulting in less stable interference nulling.
@@
Computational Complexity: The computational complexity of Algorithm 1 is dominated by computing the projections inside the loop. For instance, in (21), its main part, i.e., $\bar{\mathbf{A}}^{\ast}\left(\bar{\mathbf{A}}^{T}\bar{\mathbf{A}}^{\ast}\right)^{-1}\bar{\mathbf{A}}^{T}$, can be computed outside the loop as it only depends on the interference channels matrix. Hence, its computational complexity can be written as $\mathcal{O}\left(K^{6}+K^{2}N^{4}+K^{4}N^{2}+K^{2}N^{2}\right)$. In addition, the lossless and reciprocal projections inside the loop have a computational complexity of $\mathcal{O}\left(tN^{3}\right)$ and $\mathcal{O}\left(tN^{2}\right)$, where $t$ is the number of iterations. Consequently, the overall algorithm’s computational complexity is $\mathcal{O}\left(tN^{3}+tN^{2}+K^{6}+K^{2}N^{4}+K^{4}N^{2}+K^{2}N^{2}\right)$. If $N=2K-1$, the computational complexity can be written as $\mathcal{O}\left(K^{6}\right)$, otherwise if $N>>K$, then the computational complexity can be written as $\mathcal{O}\left(K^{2}N^{4}\right)$. Moreover, the Numerical Results and Discussions Section will show that initializing based on passive MRT solution, (14), improves the sum rate performance compared to the random initialization.
@@
Channel State Information Requirements: The interference channels matrix $\bar{\mathbf{A}}$ is needed to compute $\boldsymbol{\Theta}^{\star}$. This can be found by pre-designing the variation of $\mathbf{\Theta}$ for the fully-connected BD-RIS architecture to be further used in the training process [29].
IV Stage 2: BS Power Control/Precoding
@@
IV-A Power Control/Precoding for Passive MRT
Since the passive MRT presented in Sec. III-A does not set the interference terms to zero, a low-complexity precoder can be considered at the BS such as ZF. Consequently,
@@ IV-A Power Control/Precoding for Passive MRT
such that $p_{k}=\frac{P_{\max}}{K}$, $\forall k$. Hence, it is agnostic to the equivalent channels matrix. On the other hand, while the equivalent channels matrix is known at the BS, the optimized power can be designed to maximize the sum rate of the system. Therefore, the sum rate maximization problem can be formulated as [31, Eq. (3)],
@@
IV-B Power Control for Passive Interference Nulling
Since the passive interference nulling presented in Sec. III-B achieves ZF when sufficient number of REs exist (typically $N\geq 2K-1$), $\mathbf{E}$ is a diagonal matrix. Hence, the transmission results in $K$ virtual parallel streams. Therefore, full precoding is not required to maximize the sum rate. Alternatively, water-filling power allocation can be used, which is optimal for (25a). The closed-form water-filling solution is given as
@@ V-B Notations for Proposed Designs and Benchmarks
Various benchmarks are used to compare the performance of our proposed designs for the BD-RIS. The notations for the proposed designs are given as follows:
  • BD-RIS MRT denotes the passive MRT solution introduced in Sec. III-A.
  • BD-RIS Null denotes the passive interference nulling introduced in Sec. III-B. Both random and MRT initializations are considered for Algorithm 1.
The notations for the processes at the BS are given as
@@
  • RM denotes the rate maximization shown in (25a), (26).
  • ZF denotes the zero-forcing shown in (23).
The benchmarks belong to two main categories: MIMO relays and RIS. Under MIMO relays, noiseless non-regenerative MIMO relays are considered with $N$-transmit, $N$-receive antennas and $36$ dBm power budget operating in full-duplex without any self-interference. Two-stage design is considered:
  • nMIMO ZF denotes performing ZF at the MIMO relay, while uniform power is performed at the BS. The computational complexity at the BS is negligible, whereas the computational complexity at the relay is $\mathcal{O}(N^{3})$ which comes from computing the Moore-pseudoinverse of $\mathbf{G}$.
  • nMIMO MRT denotes performing MRT at the MIMO relay, while ZF precoding is performed at the BS. The computational complexity at the relay is $\mathcal{O}(N^{2})$ which comes from MRT computation based on $\mathbf{G}$.
D-RIS and BD-RIS are two RIS technologies that are considered. For D-RIS,
  • D-RIS Null denotes interference nulling for D-RIS following [10].
Also, joint design and various two-stage designs are considered for BD-RIS:
@@
  • BD-RIS Joint denotes the joint-design following [18], which has a computational complexity of $\mathcal{O}(t_{1}t_{2}N^{3})$ for $N>>K$, where $t_{1}$ and $t_{2}$ are the number of iterations for the outer and inner loops.
@@
  • BD-RIS Max-F denotes the Frobenius norm-based maximum channel gain design for the BD-RIS scattering matrix. It aims to $\max_{\boldsymbol{\Theta}}\left\|\mathbf{E}\right\|^{2}_{F}$, which has a relaxed closed-form solution that follows [28] and its symmetric unitary solution is given as $\boldsymbol{\Theta}^{\star}=\Pi_{\mathcal{S}_{\mathrm{loss}}}\left(\Pi_{\mathcal{S}_{\mathrm{recip}}}\left(\mathrm{reshape}\left(\sqrt{N}\mathbf{v}_{\max},[N,N]\right)\right)\right)$, where $\mathbf{v}_{\max}$ is the Eigenvector of $\mathbf{A}^{H}\mathbf{A}$ associated with its dominant Eigenvalue. Its computational complexity comes from computing the SVD of an $N^{2}\times N^{2}$ matrix and is $\mathcal{O}(N^{6})$.
  • BD-RIS Max-l2 denotes the $l^{2}$ norm-based maximum channel gain design for the BD-RIS scattering matrix. It aims to $\max_{\boldsymbol{\Theta}}\left\|\mathbf{E}\right\|^{2}_{2}$, which has a symmetric unitary solution that is described in [19]. Its computational complexity comes from computing the SVD of $\mathbf{H}$ and $\mathbf{W}$ and is $\mathcal{O}(N^{3})$.
  • BS ZF denotes the the specular reflection by BD-RIS, i.e., $\boldsymbol{\Theta}^{\star}=\mathbf{I}$, where the BS performs ZF based on $\mathbf{E}$ such that $\mathbf{P}=\mathbf{E}^{{\dagger}}\sqrt{\frac{P_{\max}}{\mathrm{Tr}\left(\mathbf{E}^{{\dagger}H}\mathbf{E}\right)}}$ and has a computational complexity of $\mathcal{O}(K^{3})$.
@@
  • BS MRT denotes the the specular reflection by BD-RIS, i.e., $\boldsymbol{\Theta}^{\star}=\mathbf{I}$, where the BS performs MRT based on $\mathbf{E}$ such that $\mathbf{P}=\mathbf{E}^{H}\sqrt{\frac{P_{\max}}{\mathrm{Tr}\left(\mathbf{E}^{H}\mathbf{E}\right)}}$ and has a computational complexity of $\mathcal{O}(K^{2})$.
@@
Except for nMIMO ZF and BS MRT, the rest of the above mentioned two-stage designs consider performing ZF at the BS following (23), which has a computational complexity of $\mathcal{O}(K^{3})$ that comes from computing the Moore-pseudoinverse of $\mathbf{E}$.
Figure 3: Power and power ratio for BD-RIS considering MRT and Max-F for $p_{k}=1,\forall k$ and $K=5$.

### v2

@@ III-B Passive Interference Nulling
Such problem is non-convex and challenging to solve. Therefore, AO is used to find $\boldsymbol{\Theta}$ that complies with the previously mentioned constraints by projecting a certain initial solution into $\mathcal{S}_{\mathrm{null}}$, $\mathcal{S}_{c_{1}}$ and $\mathcal{S}_{c_{2}}$ alternatively. Fortunately, the projections to $\mathcal{S}_{c_{1}}$ and $\mathcal{S}_{c_{2}}$ have been presented with closed-form solutions in (15), (19)–(20), while the interference nulling projection follows [10, Eq. (21a)] such that
$\Pi_{\mathcal{S}_{\mathrm{null}}}\left(\boldsymbol{\theta}_{c}\right)=\boldsymbol{\theta}_{c}-\bar{\mathbf{A}}_{c}^{\ast}\left(\bar{\mathbf{A}}_{c}^{T}\bar{\mathbf{A}}_{c}^{\ast}\right)^{-1}\bar{\mathbf{A}}_{c}^{T}\boldsymbol{\theta}_{c}$ (36)
Input: $\bar{\mathbf{A}}_{c},\,G,\,\mathrm{Iter},\,\epsilon,\,\varepsilon$
@@
1 Set $t=0$; Initialize $\boldsymbol{\Theta}^{(t)}$; Extract $\boldsymbol{\theta}_{c}^{(t)}$ (28)
2 if G=1 then
3       $[\tilde{\boldsymbol{\theta}}_{c}^{(t)}]_{i}=\Pi_{\mathcal{S}_{\mathrm{unim}}}\left([{\boldsymbol{\theta}}_{c}^{(t)}]_{i}\right)$, $\forall i\in\{1,2,\ldots,N\}$
4       $\tilde{\boldsymbol{\Theta}}^{(t)}=\mathrm{diag}\left(\tilde{\boldsymbol{\theta}}_{c}^{(t)}\right)$
5else
6       for $g=:1:G$ do
7             $\boldsymbol{\Theta}_{g}^{(t)}\!=\!\mathrm{unvec}\!\left([\boldsymbol{\theta}_{c}^{(t)}]_{(g-1)N_{g}^{2}+1:gN_{g}^{2}},\,[N_{g},N_{g}]\right)$
8             $\tilde{\boldsymbol{\Theta}}_{g}^{(t)}=\Pi_{\mathcal{S}_{\mathrm{sym}}}\left(\Pi_{\mathcal{S}_{\mathrm{uni}}}\left(\boldsymbol{\Theta}_{g}^{(t)}\right)\right)$
9
10      $\tilde{\boldsymbol{\theta}}_{c}^{(t)}=\left[\mathrm{vec}\left(\tilde{\boldsymbol{\Theta}}_{1}^{(t)}\right)^{T},\ldots,\mathrm{vec}\left(\tilde{\boldsymbol{\Theta}}_{G}^{(t)}\right)^{T}\right]^{T}$
11       $\tilde{\boldsymbol{\Theta}}^{(t)}=\mathrm{blkdiag}\left(\tilde{\boldsymbol{\Theta}}_{1}^{(t)},\ldots,\tilde{\boldsymbol{\Theta}}_{G}^{(t)}\right)$
12
13for $t=:1:\mathrm{Iter}$ do
14       $\boldsymbol{\theta}_{c}^{(t)}=\Pi_{\mathcal{S}_{\mathrm{null}}}\left(\tilde{\boldsymbol{\theta}}_{c}^{(t-1)}\right)$
15       Repeat Lines 222210101010
16       $\delta=\frac{\left\|\bar{\mathbf{A}}_{c}^{T}\tilde{\boldsymbol{\theta}}_{c}^{(t)}\right\|^{2}_{2}-\left\|\bar{\mathbf{A}}_{c}^{T}\tilde{\boldsymbol{\theta}}_{c}^{(t-1)}\right\|^{2}_{2}}{\left\|\bar{\mathbf{A}}_{c}^{T}\tilde{\boldsymbol{\theta}}_{c}^{(t-1)}\right\|^{2}_{2}}$
17       if $\delta\leq\varepsilon$ or ${\scriptstyle{\left\|\bar{\mathbf{A}}_{c}^{T}\tilde{\boldsymbol{\theta}}_{c}^{(t)}\right\|^{2}_{2}}}<\epsilon$ or $t>\mathrm{Iter}$ then
18             break and go to Line $17$
19
@@ 15       Repeat Lines 222210101010
Figure 2: Convergence of Algorithm 1 vs. number of iterations for passive interference nulling using fully-connected “FC”, group-connected “GC”, and single-connected “SC” BD-RISs with random initialization, where $K=8$ and $N=144$. a) Nulling condition $l^{2}$ norm. b) Nulling condition $l^{2}$ norm relative change.
TABLE I: Summary of the computational and circuit topology complexities for the proposed BD-RIS designs.
BD-RIS Design Computational Complexity Circuit Topology
Complexity
Single-Connected Passive MRT $\mathcal{O}\left(KN+N\right)$ $N$
Passive Interference Nulling $\mathcal{O}\left(tN\!+K^{6}\!+K^{4}N\!+K^{2}N^{2}\!+K^{2}N\right)$
Fully-Connected Passive MRT $\mathcal{O}\left(N^{3}+KN^{2}+N^{2}\right)$ $\frac{N(1+N)}{2}$
Passive Interference Nulling $\mathcal{O}\left(tN^{3}\!+tN^{2}\!+K^{6}\!+K^{4}N^{2}\!+K^{2}N^{4}\!+K^{2}N^{2}\right)$
Group-Connected Passive MRT $\mathcal{O}\left(GN_{g}^{3}+GKN_{g}^{2}+GN_{g}^{2}\right)$ $\frac{N(1+N_{g})}{2}$
Passive Interference Nulling $\mathcal{O}\left(tGN_{g}^{3}\!+tGN_{g}^{2}\!+K^{6}\!+K^{4}GN_{g}^{2}\!+K^{2}G^{2}N_{g}^{4}\!+K^{2}GN_{g}^{2}\right)$
@@
A pseudo-code that illustrates the steps to obtain $\boldsymbol{\Theta}^{\star}$ through AO is given in Algorithm 1. The algorithm needs the interference channels matrix $\bar{\mathbf{A}}_{c}$ and the number of groups $G$ as inputs. It starts by initializing the scattering matrix, which can be either a random initialization or a specific initialization. In addition, the initialized scattering matrix is projected to $\mathcal{S}_{c_{1}}$ and $\mathcal{S}_{c_{2}}$ in Lines 222210101010. Lines 1111111116161616 illustrate the iterative AO which projects the $\boldsymbol{\Theta}$ from the previous iteration to interference nulling set and $\mathcal{S}_{c_{1}}$ and $\mathcal{S}_{c_{2}}$. This continues for a certain number of iterations until the breaking condition on line $15$ is met. On line $14$, $\delta$ is computed which is the $l^{2}$ norm for the nulling condition difference between the current iteration and the previous iteration. Finally, $\boldsymbol{\Theta}^{\star}$ that achieves interference nulling and belongs to $\mathcal{S}_{c_{1}}$ and $\mathcal{S}_{c_{2}}$ is returned. It is worth noting that $\epsilon$ and $\varepsilon$ determine the algorithm convergence speed, where the former is the relative rate of change tolerance and the latter is the nulling condition norm tolerance.
@@
Fig. 2 compares the convergence of Algorithm 1 for fully-connected, group-connected and single-connected BD-RISs. Random initialization is assumed for both, with $N=144$, $K=8$, $\epsilon=10^{-10}$, $\varepsilon=10^{-6}$, $\mathrm{Iter}=10^{4}$, and $100$ realizations are considered for Monte-Carlo simulation. Also, i.i.d Rayleigh fading channels are considered. The curves represent the average convergence of the algorithm over $10^{4}$ iterations, while the shades represent the possible variations, i.e., the maximum and minimum convergence. It can be observed that fully-connected BD-RIS achieves a rapid interference nulling, reaching $10^{-8}$ with significantly fewer iterations compared to single-connected BD-RIS. In addition, the relative change for the fully-connected BD-RIS stabilizes after fewer iterations compared to the single-connected BD-RIS. Moreover, the the fully-connected BD-RIS maintains a stable pattern, effectively achieving interference nulling, unlike the single-connected BD-RIS which is more dependent on the initialization of the scattering matrix, resulting in less stable interference nulling. Furthermore, the behavior of the group-connected BD-RIS falls between the two extremes. Specifically, as the group size approaches $N$, the performance becomes closer to that of a fully-connected BD-RIS. Conversely, as the group size approaches $1$, the performance becomes similar to that of the single-connected BD-RIS.
@@
Computational Complexity: The computational complexity of Algorithm 1 is dominated by computing the projections inside the loop. For instance, in (36), its main part, i.e., $\bar{\mathbf{A}}_{c}^{\ast}\left(\bar{\mathbf{A}}_{c}^{T}\bar{\mathbf{A}}_{c}^{\ast}\right)^{-1}\bar{\mathbf{A}}_{c}^{T}$, can be computed outside the loop as it only depends on the interference channels matrix. Hence, its computational complexity can be written as $\mathcal{O}\left(K^{6}+K^{4}GN_{g}^{2}+K^{2}G^{2}N_{g}^{4}+K^{2}GN_{g}^{2}\right)$. For the single-connected architecture, the unit-modulus projection has a computational complexity of $\mathcal{O}\left(tN\right)$, while the symmetric and unitary projections associated with group-connected architecture have a computational complexity of $\mathcal{O}\left(tGN_{g}^{3}\right)$ and $\mathcal{O}\left(tGN_{g}^{2}\right)$. Consequently, the overall algorithm’s computational complexity is $\mathcal{O}\left(tGN_{g}^{3}\!+tGN_{g}^{2}\!+K^{6}\!+K^{4}GN_{g}^{2}\!+K^{2}G^{2}N_{g}^{4}\!+K^{2}GN_{g}^{2}\right)$. Typically, when $N\gg K$ the computational complexity can be written as $\mathcal{O}\left(K^{2}G^{2}N_{g}^{4}\right)$. Moreover, the Numerical Results and Discussions Section will show that initializing based on passive MRT solution, (14), (23), improves the sum rate performance compared to the random initialization.
@@
Channel State Information Requirements: The interference channels matrix $\bar{\mathbf{A}}_{c}$ is needed to compute $\boldsymbol{\Theta}^{\star}$. This can be found by pre-designing the variation of $\mathbf{\Theta}$ for the single/group-connected BD-RIS architecture to be further used in the training process [35].
For the readers convenience, Table I summarizes the computational complexity and circuit toplogy complexity for the different BD-RIS passive MRT and passive interference nulling designs.
IV Stage 2: BS Power Allocation/Precoding
@@
IV-A Power Allocation/Precoding for Passive MRT
Since the passive MRT presented in Sec. III-III-A does not set the interference terms to zero, a low-complexity precoder can be considered at the BS such as ZF. Consequently,
@@ IV-A Power Allocation/Precoding for Passive MRT
such that $p_{k}=\frac{P_{\max}}{K}$, $\forall k$. Hence, it is agnostic to the equivalent channels matrix. On the other hand, while the equivalent channels matrix is known at the BS, the optimized power can be designed to maximize the sum rate of the system. Therefore, the sum rate maximization problem can be formulated as [37, Eq. (3)],
@@
IV-B Power Allocation for Passive Interference Nulling
Since the passive interference nulling presented in Sec. III-III-B achieves ZF when sufficient number of REs exist, i.e., $N\geq\left\lceil\frac{4K(K-1)}{1+N_{g}}\right\rceil$, $\mathbf{E}$ is a diagonal matrix. Hence, the transmission results in $K$ virtual parallel streams. Therefore, full precoding is not required to maximize the sum rate. Alternatively, water-filling power allocation can be used, which is optimal for (40a). The closed-form water-filling solution is given as
@@ V-B Notations for Proposed Designs and Benchmarks
Various benchmarks are used to compare the performance of our proposed designs for the BD-RIS. The notations for the stage 1 proposed designs are given as follows:
  • MRT denotes the passive MRT solution introduced in Sec. III-III-A.
  • Null denotes the passive interference nulling introduced in Sec. III-III-B, where MRT initializations (14), (23) are considered for Algorithm 1.
The notations for the stage 2 processes at the BS are:
  • ZF denotes the zero-forcing precoder shown in (38).
@@
  • RM denotes the rate maximization optimization $(P5)$ shown in (40a)–(40c) and the closed-form water-filling shown in (41).
The considered benchmarks are:
@@
  • Joint denotes the joint-design for the BD-RIS following [17], which has a computational complexity of $\mathcal{O}(t_{1}t_{2}t_{3}GN_{g}^{3})$ for the group-connected, where $t_{1}$, $t_{2}$ and $t_{3}$ are the number of iterations for the algorithm loops.
@@
  • Max-F denotes the Frobenius norm-based maximum channel gain design for the BD-RIS scattering matrix. It aims to $\max_{\boldsymbol{\Theta}}\left\|\mathbf{E}\right\|^{2}_{F}$, which has a relaxed closed-form solution that follows [34]. The optimal solution that belongs to $\mathcal{S}_{\mathrm{SC}_{1}}$ and $\mathcal{S}_{\mathrm{SC}_{2}}$ for the single-connected is given as
    $\left[\boldsymbol{\Theta}^{\star}\right]_{i,j}=\left\{\begin{array}[]{lr}\Pi_{\mathcal{S}_{\mathrm{unim}}}\left(\left[\mathbf{v}_{\max}\right]_{i}\right),&i=j\\ 0,&i\neq j\end{array}\right.$ (42)
    where $\mathbf{v}_{\max}$ is the Eigenvector of $\mathbf{A}_{c}^{H}\mathbf{A}_{c}$ associated with its dominant Eigenvalue. The optimal solution that belongs to $\mathcal{S}_{\mathrm{GC}_{1}}$ and $\mathcal{S}_{\mathrm{GC}_{2}}$ for the group-connected is given as (22) and
@@
    $\boldsymbol{\Theta}_{g}^{\star}=\Pi_{\mathcal{S}_{\mathrm{symuni}}}\left(\!\mathrm{unvec}\left(\!\sqrt{N_{g}}\mathbf{v}_{g,\max},[N_{g},N_{g}]\!\right)\!\right)$ (43)
@@
    where $\mathbf{v}_{g,\max}=\left[\mathbf{v}_{\max}\right]_{(g-1)N_{g}^{2}+1:gN_{g}^{2}}$. Its computational complexity comes from computing SVD for $GN_{g}^{2}\times GN_{g}^{2}$ matrix and is $\mathcal{O}(G^{3}N_{g}^{6})$. At the BS, ZF is performed following (38), which has a computational complexity of $\mathcal{O}(K^{3})$ that comes from computing the Moore-pseudoinverse of $\mathbf{E}$.
  • Spec-ZF denotes the the specular reflection by BD-RIS, i.e., $\boldsymbol{\Theta}^{\star}=\mathbf{I}$, where the BS performs ZF based on $\mathbf{E}$ such that $\mathbf{P}=\mathbf{E}^{{\dagger}}\sqrt{\frac{P_{\max}}{\mathrm{Tr}\left(\mathbf{E}^{{\dagger}H}\mathbf{E}\right)}}$ and has a computational complexity of $\mathcal{O}(K^{3})$.
  • Spec-MRT denotes the the specular reflection by BD-RIS, i.e., $\boldsymbol{\Theta}^{\star}=\mathbf{I}$, where the BS performs MRT based on $\mathbf{E}$ such that $\mathbf{P}=\mathbf{E}^{H}\sqrt{\frac{P_{\max}}{\mathrm{Tr}\left(\mathbf{E}^{H}\mathbf{E}\right)}}$ and has a computational complexity of $\mathcal{O}(K^{2})$.
Figure 3: Desired and interference powers for fully-connected “FC” and group-connected “GC” BD-RIS considering MRT and Max-F for $p_{k}=1,\forall k$, $K=5$ and $N=16$.
Figure 4: Sum rate for passive interference nulling using fully-connected “FC”, group-connected “GC”, and single-connected “SC” BD-RISs with MRT initialization, where a) x-axis is transmit power and $K=5$. b) x-axis is number of users and $P_{\max}=5$ dBm.
(a) Passive MRT
(b) Passive Interference Nulling
Figure 5: Sum rate vs. transmit power for the stage 2 processes at the BS using fully-connected “FC”, group-connected “GC”, and single-connected “SC” BD-RISs, where $K=5$, $P_{\max}=5$ dBm, $N=64$. a) Passive MRT. b) Passive interference nulling.